The length of a rectangle is 4cm more than twice its width. The perimeter is 86cm. Find its width and length. Thank you so much.
Answer by mananth(16949)

Perimeter = 86
Length 1 ( 2 x + 4 )
Width = x ft

Perimeter = 2((2x+4 )+1x)
Perimeter = 2(2x+4+ 1x)
Perimeter = 4x+8+2x
86 =6x+8
86-8= 6x
78=6 x
x=13 cm Width
Length =30 cm